OpenStudy (anonymous):

What is the slope intercept form that passes through the given point and is perpendicular (-8,5),y=-4x+2? What is the answer?! Help!! :)

OpenStudy (marigirl):

right, so we are after an equation that is perpendicular to y=-4x+2 isnt it?

OpenStudy (marigirl):

m1*m2=-1 .. we can find the gradient of the perpendicular line through this equation. we know m1=-4 so then -4*m2=-1. what would be m2 (or the gradient of the perpendicular line)

OpenStudy (anonymous):

But what would the answer be?

OpenStudy (marigirl):

i am working through the ans with you. first we need to gradient of the perpendicular line. which is my prev post. would u be able to tell me what m2 is.

OpenStudy (marigirl):

can u see that using the equation

OpenStudy (marigirl):

m1*m2=-1 -4*m2=-1 therefore, m2 must be (1/4) are u with me on that

OpenStudy (anonymous):

Y=1/4+x answer?

OpenStudy (marigirl):

y=(1/4)x+c is the equation. now we need to use the coordinate( -8,5) to compete the equation and find c

OpenStudy (marigirl):

using the coordinate, i am going to substitute in the x and y values to solve for c 5=(1/4)(-8)+c now we solve for c, which is the y intercept or constant 5=(-8/4)+c 5=-2+c c=7 y=(1/4)x+7